Write an equation in point-slope form for the line parallel to -18x + 2y = 7 that contains (3, 1).
Answer by linht117(37)

point-slope= y - y1 = (m) (x - x1)
line parallel= same slope
line perpendicular= opposite reciprocal slope
put -18x+2y=7 in y-intercept form
-18x + 18x + 2y = 7 + 18x
2y/2 = (7 + 18x)/2
y = 9x + 7/2
slope is 9
now use (3,1)& (9) in point-slope
y - 1 = (9) (x - 3)